Estimates of a firm's returns revealed that β of the firm was 1.1, and the historical average market return over the past 5 years has been 13.1%, and the current risk-free rate is 2.84%. What is the expected return of this stock based on the Capital Asset Pricing Model (CAPM). The firm just recently released a dividend for $0.99 per share, given an expected dividend growth rate of 13.60%, what is the price per share of the stock?
Answer #1

expected return=risk-free rate +Beta*(market rate- risk-free rate )

=2.84+1.1*(13.1-2.84)

=14.126%

Current price=D1/(Required return-Growth rate)

=(0.99*1.136)/(0.14126-0.136)

=$213.81(Approx).